Grade 12th passPhysical Chemistry

Enthalpy change for vapourization of liquid a at 200k and 1 atm is 22kj/mol. Find change in entropy at 200k .the normal bp is 300k . Cp ,m rxn (-10 j/mol)

haroon
11 Years agoGrade 12th pass

1 Answer

MANOJ
11 Years ago
Since phase change reactions are equilibrium reactions dG=0;
Also dG=dH-TdS
dH=TdS
dS=22000/200=110J/molK